OpenStudy (anonymous):

prove (u-v)dot(u+v) = |u|^2-|v|^2

OpenStudy (anonymous):

(u-v)(u+v) = u^2 +uv-uv-v^2 = u^2 - v^2 |u|^2 = has no meaning cause power (2) always positive

OpenStudy (anonymous):

no this is vector

OpenStudy (anonymous):

ahhh linear algebra